At a fast food restaurant, one burger, one order of fries, and one drink cost $5; three burgers, two orders of fries, and two drinks cost $12.50; and two burgers, four orders of fries, and three drinks cost $14. Find the cost of one burger.
Answer by amarjeeth123(569)

Let the cost of one burger be x.
Let the cost of one order of fries be y.
Let the cost of one drink be z.
x+y+z=5...........equation 1
3x+2y+2z=12.50....equation 2
2x+4y+3z=14.......equation 3
Multiplying equation 1 by 2 we get
2x+2y+2z=10.......equation 4
Subtracting equation 4 from equation 2 we get,
3x+2y+2z-2x-2y-2z=12.5-10
x=2.5
The cost of one burger is $2.50.